Consider an annuity of payments of $1000 at the end of every second year. What is the present value of this annuity if it runs for ten years and the interest rate is 7%?
1
Expert's answer
2020-10-05T18:39:56-0400

PV of Ordinary Annuity = C * {1-(1+i)-n/i}


Effective interest rate at the end of every 2 years= (1+0.07)2-1= 0.14


Substitute the values in the formula C= $ 10,000

i= 14%

n= 10/2= 5


PV = $10,000 * { 1-(1+0.14)-5/ 0.14}


= $10,000 * 3.432


= $34,320